OpenStudy (anonymous):

Identify which technique would be best applied to factor the polynomial below. 2x^2+xy+6x+3y A. difference of two squares B. grouping C. greatest common factor D. perfect square trinomial also with x^2-25 as well. ( i think its d for that one)

OpenStudy (xapproachesinfinity):

why do you think D?

OpenStudy (anonymous):

because 25 is a perfect square idk honestly just a guess

OpenStudy (xapproachesinfinity):

this is straight forward \[x^2-25=x^2-5^2\]

OpenStudy (xapproachesinfinity):

so that's a difference of two perfect square x^2 and 5^2

OpenStudy (anonymous):

oh okay thank you!!! any idea about the first one?

OpenStudy (xapproachesinfinity):

the first one would be like this \[2x^2+xy+6x+3y=2x^2+6x+xy+3y=2x(x+3)+y(x+3)\\=(x+3)(2x+y)\]

OpenStudy (xapproachesinfinity):

that called grouping

OpenStudy (anonymous):

thank you so much

OpenStudy (xapproachesinfinity):

weclome